OpenStudy (anonymous):

The French Club is sponsoring a bake sale. If its goal is to raise at least $140, how many pastries must it sell at $3.50 to meet that goal? Write and solve an inequality.

OpenStudy (anonymous):

The French Club is sponsoring a bake sale. If its goal is to raise at least $140, how many pastries must it sell at $3.50 to meet that goal? Write and solve an inequality. 3.50p is equal to or greater than 140; p is equal to or greater than 490 140p is equal to or greater than 3.50; p is equal to or greater than 40 3.50p is equal to or greater than 140; p is equal to or greater than 136.5 3.50p is equal to or greater than 140; p is equal to or greater than 40

OpenStudy (lifeisadangerousgame):

p is the variable for pastries right?

OpenStudy (anonymous):

Right @LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

I'm going to write out each problem so that it's easier to read visually. \[3.50p \ge 140; p \ge 490\] \[1.40p \ge 3.50; p \ge 40\] \[3.50p \ge 140; p \ge 136.5\] \[3.50p \ge 140; p \ge 40\] Are those right?

OpenStudy (lifeisadangerousgame):

While you're doing that, I'll show you how to solve it. Your goal is to sell enough pastries that you make 140 dollars or more. So what you do is Least amount you want to make/price per pastry --> 140/3.50 = 40. So 40 * 3.50 equals 140, which is the minimum amount of pastries you have to sell in order to make the minimum amount of money you want to make. Are you following so far? Any idea of what the answer is based on what we just did?

OpenStudy (anonymous):

Answer #2 is 140, not 1.40...Sorry if that confused you

OpenStudy (anonymous):

140p, not 1.40p

OpenStudy (lifeisadangerousgame):

Oh oops, I apologize. \[140p \ge 3.50; p \ge 40\] Instead Now, are you following along or what's happening? c:

OpenStudy (anonymous):

and from what you just typed, I may have misunderstood, isnt it B? 140p (is equal to or greater than) 3.50;p (is equal to or greater than) 40?

OpenStudy (anonymous):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

Unfortunately no, because 140p means 140 dollars times the amount of pastries you sell. You are selling pastries for 3.50, not 140 (thank goodness!). Should I explain it a bit better?

OpenStudy (anonymous):

Please. Sorry. Kinda confused

OpenStudy (anonymous):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

It's totally fine. So what you're trying to find out is how many pastries you need to sell, for 3.50 each, in order to make it to the minimum amount of money you need to make, which is $140. p stands for pastries, or the amount of pastries you sell. So you could either plug in numbers for p, until you got to 40 (which is $3.50 * 40 and equals 140) or divide 140 by 3.50 to get 40. 40 pastries is the amount of pastries you have to sell to make at LEAST $140. Following so far?

OpenStudy (anonymous):

Yeah.

OpenStudy (lifeisadangerousgame):

Okay so here's what we know. The minimum amount of money we need to make is: $140 The minimum amount of pastries we need to sell is: 40 pastries We need to make $140 dollars OR GREATER and we need to sell 40 pastries OR MORE. Making more sense now?

OpenStudy (anonymous):

Yeah. It's starting to. (thanks for walking me through this, by the way)

OpenStudy (lifeisadangerousgame):

Of course! I want you to learn this! So now, based on this: We need to make $140 dollars OR GREATER and we need to sell 40 pastries OR MORE. We now know that the answer is is D. Because it matches the sentence above, it says the exact same thing as the sentence, but in numbers. "You must make 3.50*p and it has to be more than or equal to $140 and p must be more than or equal to 40 pastries." Making sense?

OpenStudy (anonymous):

lol i rewarded you for excellent explanation

OpenStudy (lifeisadangerousgame):

Haha thank you Student xD This is helping me with remembering this stuff

OpenStudy (anonymous):

Yeah it makes sense....Thank you so much.

OpenStudy (lifeisadangerousgame):

yay!! Definitely do more of these so that you really get it in your brain. Congratulations, you can now math:D

OpenStudy (anonymous):

check your problem life i hope it may help lol

OpenStudy (lifeisadangerousgame):

Will do! Gonna finish up this other person's stuff then I'll head on over. Thanks for replying!